A train start from rest accelerates uniformly so that it attain a speed of 400-1 in 20s. It Travel at this speed for 40s and is then brought to rest with uniform retardation after 30s




What is the



The total distance covered



The retardation of the train

1
Expert's answer
2022-03-20T18:48:11-0400

The total distance traveled by a train

"d=d_1+d_2+d_3"

"d=\frac{v+0}{2}t_1+vt_2+\frac{0+v}{2}t_3"

"d=\frac{400}{2}\times20+400\times 40+\frac{400}{2}\times 30\\
=26000\:\rm m"

The retardation of the train

"a_3=\frac{\Delta v_3}{\Delta t_3}=\frac{-400}{30}=-13.3\;\rm m/s^2"